# EKF–SINDy: Empowering the extended Kalman filter with sparse identification of nonlinear dynamics

### v1

@@ 2.2 State estimation of autonomous dynamical system
The mapping $\mathbf{f}$ is only sometimes known, but, even if this is the case, the computation of $\mathbf{F}$ can be rather involved, while its numerical calculation can be computationally expensive. These reasons suggest to use SINDy to replace $\mathbf{f}$. Indeed, by expressing $\mathbf{f}$ as the linear combination of a set of predetermined functions collected in a library, as SINDy does, its (partial) derivatives can be simply obtained by combining the (partial) derivatives of the library functions. Analogously, even though $\mathbf{h}$ does not describe a dynamic system, it could be also identified using SINDy, thus greatly simplifying the computation of $\mathbf{H}$. However, in general the function $\mathbf{h}$ is directly set by means of more straightforward techniques. In the first case study proposed in this work, system state and observations represent the same dynamic quantities, and they do not require a distinct identification for $\mathbf{h}$. In the second case study, $\mathbf{h}$ has been determined from the time-delay embedding required by SINDy to (fully) learn $\mathbf{f}$.
For sake of generality, we now consider to use SINDy to model both $\mathbf{f}$ and $\mathbf{h}$. A detailed procedure is reported for $\mathbf{f}$, but the same strategy can also be applied to $\mathbf{h}$. The discussion is related to autonomous systems, i.e., systems without external forcing; in the next section, this assumption will be relaxed to allow for non-autonomous systems.
@@
Similarly, a matrix $\dot{\mathbf{X}}$ is constructed by collecting the time derivatives $\dot{\mathbf{x}}$. We remark that the number of snapshots should not be necessarily equal to $T$. If $\mathbf{f}$ is available, Eq. (1) can be used to generate $\mathbf{X}$ and $\dot{\mathbf{X}}$. Otherwise, $\mathbf{f}$ can be identified from experimental data exploiting noise tolerant versions of SINDy, see, e.g., art:Messenger21 ; art:Fasel22 ; art:Hirsh22 ; art:Gao23 .
A library $\boldsymbol{\Theta}_{f}(\mathbf{x})=\left[\theta_{1}(\mathbf{x}),\ldots,\theta_{p}(\mathbf{x})\right]\in\mathbb{R}^{p}$ of $p$ candidate functions to describe the dynamics of the data is selected, and the matrix $\boldsymbol{\Theta}_{f}(\mathbf{X})\in\mathbb{R}^{T\times p}$ is constructed by applying $\boldsymbol{\Theta}$ to the rows of $\mathbf{X}$, e.g., as in the following:
@@
It is worth noting that any type of function can be utilised to form the function library. For instance, in Eq. (10) we have employed constant, polynomial and trigonometric functions. Including polynomial functions is effective in identifying dominant dynamic behaviour, owing to the Taylor expansion of the function governing the system dynamics art:Champion19 . The way in which the quadratic nonlinearities $\mathbf{X}^{P_{2}}$ are expressed is now explicitly detailed:
@@
The function $\mathbf{f}$ can be then expressed as the linear combination of some of these candidate functions. The weighting coefficients of the combination are stored in the matrix $\boldsymbol{\Xi}_{f}=\left[\boldsymbol{\xi}^{f}_{1},\boldsymbol{\xi}^{f}_{2},\ldots,\boldsymbol{\xi}^{f}_{n}\right]$ with $\boldsymbol{\xi}^{f}_{i}\in\mathbb{R}^{p}$ for $i=1,\ldots,n$. In matrix form, such combination can be rewritten as:
@@
where the equality sign holds as $\boldsymbol{\Theta}_{f}$ admits an arbitrary large number of terms, thus allowing – in principle – to perfectly reconstruct the system dynamics. As only few terms of the function library are expected to be able to describe the dynamics of the system of interest, it is assumed that $\mathbf{f}$ admits a sparse representation in $\boldsymbol{\Theta}_{f}$. Consequently, a $L^{1}$ (sparsity promoting) regularisation term is added in the least square regression used to determine the weighting coefficients, according to:
$\boldsymbol{\xi}^{f}_{i}=\underset{\boldsymbol{\xi}^{\prime}_{i}}{\arg\min}\|\dot{\mathbf{X}}_{i}-\boldsymbol{\Theta}_{f}(\mathbf{X})\boldsymbol{\xi}^{\prime}_{i}\|_{2}+L\|\boldsymbol{\xi}^{\prime}_{i}\|,$ (13)
where $L$ is a suitable regularisation parameter.
As previously mentioned, SINDy can be used to identify both $\mathbf{f}$ and $\mathbf{h}$, even though $\mathbf{h}$ does not describe a system dynamics. To cope with that, a matrix $\mathbf{Y}$ collecting snapshots of $\mathbf{y}$ is constructed in addition to $\mathbf{X}$. Finally, Eq. (12) is rewritten as:
@@
where the employed function library $\boldsymbol{\Theta}_{h}$ can be in general different from $\boldsymbol{\Theta}_{f}$. The weighting coefficient of the combination $\boldsymbol{\xi}^{h}_{i}$, stored in the matrix $\boldsymbol{\Xi}_{h}=\left[\boldsymbol{\xi}^{h}_{1},\boldsymbol{\xi}^{h}_{2},\ldots,\boldsymbol{\xi}^{h}_{o}\right]$, are determined through a least square regression similar to what done in Eq. (13), provided the substitution of $\dot{\mathbf{X}}_{i}$ by $\mathbf{Y}_{i}$, and of $\boldsymbol{\Theta}_{f}$ by $\boldsymbol{\Theta}_{h}$.
Modelling $\mathbf{f}$ and $\mathbf{h}$ with SINDy enables to easily compute the Jacobian matrices $\mathbf{F}$ and $\mathbf{H}$, as it implies to calculate the derivatives of the functions of the library $\boldsymbol{\Theta}_{f}$ and $\boldsymbol{\Theta}_{h}$. Specifically, by combining Eqs. (5) and 8 with Eqs. (12) and (14), $\mathbf{F}$ and $\mathbf{H}$ assume the following forms:
@@
2.3 Joint estimation of non-autonomous dynamic systems
@@
The goal of joint estimation is to simultaneously estimate $\mathbf{x}$ and the parameters $\boldsymbol{\phi}$ art:Mariani05 . To reach the scope, we increase the number of state variables by augmenting the state vector as $\boldsymbol{\varkappa}=[\mathbf{x}^{\top},\boldsymbol{\phi}^{\top}]^{\top}$. We consider the case in which it is impossible to associate a dynamic description to the evolution of $\boldsymbol{\phi}$. In this setting, the prediction stage of $\boldsymbol{\phi}$ is performed by a random walk driven by a process noise $\mathbf{q}_{\phi}$ as in the following:
@@
where $\mathbf{q}_{\phi}$ is assumed to be sampled from a white, zero mean stochastic process $\mathbf{q}_{\phi}\sim\mathcal{N}(\mathbf{0},\mathbf{Q}_{\phi})$ with a diagonal covariance matrix $\mathbf{Q}_{\phi}\in\mathbb{R}^{l\times l}$. Keeping the definition of $\mathbf{q}_{x}$ and $\mathbf{Q}_{x}$, we introduce a unique process noise vector $\mathbf{q}=[\mathbf{q}_{x},\mathbf{q}_{\phi}]^{\top}$ and a unique diagonal covariance matrix $\mathbf{Q}$ collecting the entries of $\mathbf{Q}_{x}$ and $\mathbf{Q}_{\phi}$.
@@
Also in case of joint estimation, it is possible – and convenient – to rely on SINDy. Specifically, the dependence of $\mathbf{f}$ on $\boldsymbol{\phi}$ can be accounted for by stacking snapshots $\boldsymbol{\varkappa}$ accounting for $I$ different realisations of $\boldsymbol{\phi}$, thereby obtaining a modified version $\mathbf{X}_{\varkappa}$ of the snapshot matrix. Thus, SINDy is employed to reconstruct the functional relation between $\mathbf{X}_{\varkappa}$ and $\dot{\mathbf{X}}$. It is unnecessary to modify the definition of $\dot{\mathbf{X}}$, as we have assumed to drive the dynamics of $\boldsymbol{\phi}$ with a random walk. Specifically, $\dot{\boldsymbol{\varkappa}}=[\dot{\mathbf{x}}^{\top},\mathbf{0}^{\top}]^{\top}$, where $\mathbf{0}\in\mathbb{R}^{l}$ is the null vector. Notably, setting $\dot{\boldsymbol{\phi}}=0$ implies that the parameter values can be updated only in the correction phase of the EKF. When used to model $\mathbf{h}$ for joint estimation, SINDy instead reconstructs the mapping between $\mathbf{X}_{\varkappa}$ and $\mathbf{Y}$. Also the definition of $\mathbf{Y}$ must not be modified as system parameters can not, in general, be observed. We highlight that the procedure can not explicitly handle the dependence on other parameters not included in $\boldsymbol{\phi}$, for which it is impossible to provide a separate estimate. This issue will be addressed in a future extension of the current work.
@@
$\dot{\mathbf{x}}(t)$ $=\mathbf{f}(\mathbf{x}(t),\mathbf{b}),$ (17a)
@@
Similar modifications apply to Eqs. (3a) and (3b). The other equations of the filter remain unchanged if we assume to know $\mathbf{b}$ that is in the case of an exogenous input. This is true, for instance, if we consider a seismic excitation thanks to the presence of monitoring networks art:DAlessandro19 . If $\mathbf{b}$ was unknown, it would be possible to modify the formulation of the filter to perform input–state–parameter estimation art:EftekharAzam15 ; art:Dertimanis19 ; art:Castiglione20 ; art:EbrahimzadehHassanabadi23 . Also this further extension of the method will be considered in future.
The external forcing term can be handled via SINDy by constructing a snapshot matrix $\mathbf{X}_{\varkappa b}$ assembling different snapshots $[\boldsymbol{\varkappa}(t_{j}),\mathbf{b}(t_{j})]^{T}$ determined for several realisations of $\boldsymbol{\phi}$. As a result, SINDy is employed to reconstruct the functional relation between $\mathbf{X}_{\varkappa b}$ and $\dot{\mathbf{X}}$ (and between $\mathbf{X}_{\varkappa b}$ and $\mathbf{Y}$ when used to model $\mathbf{h}$).
The complete procedure for the joint estimation of non-autonomous dynamic system are reported in Algorithms 1 and 2. In the algorithms, we consider the most general case in which both $\mathbf{f}$ and $\mathbf{h}$ are identified by using SINDy. The training of the SINDy model, and therefore the assembling of the snapshot matrices, must be done in an offline phase a priori with respect to the application of the procedure. Algorithm 1 is devoted to the illustration of this phase assuming to know $\mathbf{f}$ and $\mathbf{h}$. This case is of interest since it will be considered in the results section. As previously remarked, it is also possible to estimate $\mathbf{f}$ and $\mathbf{h}$ using noise tolerant versions of SINDy. During the online phase, the procedure is applied to update the estimates of the system by using incoming data. Algorithm 2 refers to this second part.
@@
12:Compute $\boldsymbol{\xi}^{f}_{i}=\underset{\boldsymbol{\xi}^{\prime}_{i}}{\arg\min}\|\dot{\mathbf{X}}_{i}-\boldsymbol{\Theta}_{f}(\mathbf{X}_{\varkappa b})\boldsymbol{\xi}^{\prime}_{i}\|_{2}+\lambda\|\boldsymbol{\xi}^{\prime}_{i}\|$ for $i=1,\ldots,n$
13:Compute $\boldsymbol{\xi}^{h}_{i}=\underset{\boldsymbol{\xi}^{\prime}_{i}}{\arg\min}\|{\mathbf{Y}}_{i}-\boldsymbol{\Theta}_{h}(\mathbf{X}_{\varkappa b})\boldsymbol{\xi}^{\prime}_{i}\|_{2}+\lambda\|\boldsymbol{\xi}^{\prime}_{i}\|$ for $i=1,\ldots,o$
@@
Algorithm 1 EKF-SINDy joint estimation ($\mathbf{x}$ and $\boldsymbol{\phi}$) of non autonomous dynamical system: offline phase ($\mathbf{f}$ and $\mathbf{h}$ available).

### v2

@@ 2.2 State estimation of autonomous dynamical system
In this subsection we present the main tools we combine with EKF to devise our proposed EKF-SINDy strategy, namely, a SINDy model and time-delay embedding. We thus explain how this combination can be exploited to perform the joint estimation of non-autonomous dynamical systems.
2.2.1 Construction of the SINDy model
The mapping $\mathbf{f}$ is only sometimes known, but, even if this is the case, the computation of $\mathbf{F}$ can be rather involved, while its numerical calculation can be computationally expensive. These reasons suggest to use SINDy to replace $\mathbf{f}$. Indeed, by expressing $\mathbf{f}$ as the linear combination of a set of predetermined functions collected in a library, as SINDy does, its (partial) derivatives can be simply obtained by combining the (partial) derivatives of the library functions. Analogously, even though $\mathbf{h}$ does not describe a dynamical system, it could be also identified using SINDy, thus greatly simplifying the computation of $\mathbf{H}$. However, in general the function $\mathbf{h}$ is directly set by means of more straightforward techniques. In the first case study proposed in this work, system state and observations represent the same dynamic quantities, and they do not require a distinct identification for $\mathbf{h}$. In the second case study, $\mathbf{h}$ has been determined from the time-delay embedding required by SINDy to (fully) learn $\mathbf{f}$.
For the sake of generality, we now consider using SINDy to model both $\mathbf{f}$ and $\mathbf{h}$. The training of SINDy is performed in an offline phase, prior to the online assimilation of incoming data performed through the EKF prediction-correction scheme. A detailed procedure is reported for $\mathbf{f}$, but the same strategy can also be applied to $\mathbf{h}$. The discussion is related to autonomous systems, i.e., systems without external forcing; in the next section, this assumption will be relaxed to allow for non-autonomous systems.
@@ 2.2.1 Construction of the SINDy model
Similarly, a matrix $\dot{\mathbf{X}}$ is constructed by collecting the time derivatives $\dot{\mathbf{x}}$. If $\mathbf{f}$ is available, Eq. (1) can be used to generate $\mathbf{X}$ and $\dot{\mathbf{X}}$. Otherwise, $\mathbf{f}$ can be identified from experimental data exploiting noise tolerant versions of SINDy, see, e.g., [34, 35, 36, 37, 38].
A library $\boldsymbol{\Theta}_{f}(\mathbf{x})=\left[\theta_{1}(\mathbf{x}),\ldots,\theta_{p}(\mathbf{x})\right]\in\mathbb{R}^{p}$ of $p$ candidate functions to describe the dynamics of the data is selected, and the matrix $\boldsymbol{\Theta}_{f}(\mathbf{X})\in\mathbb{R}^{T\times p}$ is constructed by applying $\boldsymbol{\Theta}_{f}$ to the rows of $\mathbf{X}$, e.g., as in the following:
@@
It is worth noting that any type of function can be utilised to form the function library. For instance, in Eq. (10) we have employed constant, polynomial and trigonometric functions. Including polynomial functions is effective in identifying dominant dynamic behaviour, owing to the Taylor expansion of the function governing the system dynamics [27]. The way in which the quadratic nonlinearities $\mathbf{X}^{P_{2}}$ are expressed is now explicitly detailed:
@@
The function $\mathbf{f}$ can be then expressed as the linear combination of these candidate functions. The weighting coefficients of the combination are stored in the matrix $\boldsymbol{\Xi}_{f}=\left[\boldsymbol{\xi}^{f}_{1},\boldsymbol{\xi}^{f}_{2},\ldots,\boldsymbol{\xi}^{f}_{n}\right]$ with $\boldsymbol{\xi}^{f}_{i}\in\mathbb{R}^{p}$ for $i=1,\ldots,n$. In matrix form, such combination can be rewritten as:
@@
where the equality sign holds as $\boldsymbol{\Theta}_{f}$ admits an arbitrary large number of terms, thus allowing – in principle – to perfectly reconstruct the system dynamics. As only few terms of the function library are expected to be able to describe the dynamics of the system of interest, it is assumed that $\mathbf{f}$ admits a sparse representation in $\boldsymbol{\Theta}_{f}$. Consequently, a sparsity promoting regulariser $\mathcal{L}(\boldsymbol{\xi}^{\prime}_{i})$ is added in the least square regression, used to determine the weighting coefficients, according to:
$\boldsymbol{\xi}^{f}_{i}=\underset{\boldsymbol{\xi}^{\prime}_{i}}{\arg\min}\|\dot{\mathbf{X}}_{i}-\boldsymbol{\Theta}_{f}(\mathbf{X})\boldsymbol{\xi}^{\prime}_{i}\|_{2}+\mathcal{L}(\boldsymbol{\xi}^{\prime}_{i}).$ (13)
For instance, problem (13) could be solved via LASSO regression [39], by considering $\mathcal{L}$ to be a sparsity-promoting $\ell_{1}-$norm, ı.e. $\mathcal{L}(\boldsymbol{\xi}^{\prime}_{i})=\delta_{l}\|\boldsymbol{\xi}^{\prime}_{i}\|_{1}$, where $\delta_{l}$ is the parameter setting the strength of the LASSO regularisation. Alternatively, the Sequential Thresholded Least SQuares (STLSQ) algorithm [5] could be used. It consists in iteratively solving (13) with ridge regression by using $\ell_{2}-$regularization, ı.e. $\|\boldsymbol{\xi}^{\prime}_{i}\|_{2}$, regulated by a parameter $\delta_{r}$, and setting to zero the entries of $\boldsymbol{\xi}^{\prime}_{i}$ with magnitude below a threshold $L$ at each iteration. As previously mentioned, SINDy can be used to identify both $\mathbf{f}$ and $\mathbf{h}$, even though $\mathbf{h}$ does not describe a system dynamics. To cope with that, a matrix $\mathbf{Y}$ collecting snapshots of $\mathbf{y}$ is constructed in addition to $\mathbf{X}$. Finally, Eq. (12) is rewritten as:
@@
where the employed function library $\boldsymbol{\Theta}_{h}$ can be in general different from $\boldsymbol{\Theta}_{f}$. The weighting coefficient of the combination $\boldsymbol{\xi}^{h}_{i}$, stored in the matrix $\boldsymbol{\Xi}_{h}=\left[\boldsymbol{\xi}^{h}_{1},\boldsymbol{\xi}^{h}_{2},\ldots,\boldsymbol{\xi}^{h}_{o}\right]$, are determined through a least square regression similar to what done in Eq. (13), provided the substitution of $\dot{\mathbf{X}}_{i}$ by $\mathbf{Y}_{i}$, of $\boldsymbol{\Theta}_{f}$ by $\boldsymbol{\Theta}_{h}$.
Modelling $\mathbf{f}$ and $\mathbf{h}$ with SINDy enables to easily compute the Jacobian matrices $\mathbf{F}$ and $\mathbf{H}$, as it implies to calculate the derivatives of the functions of the library $\boldsymbol{\Theta}_{f}$ and $\boldsymbol{\Theta}_{h}$. Specifically, by combining Eqs. (5) and (8) with Eqs. (12) and (14), $\mathbf{F}$ and $\mathbf{H}$ assume the following forms:
@@
2.2.2 Time-delay embedding
A possible drawback of SINDy is that its training requires knowledge of the full underlying state space variables. In many real-world applications, states could not be directly observed and some governing variables may be completely unobserved, resulting in observations ${\mathbf{y}}\in\mathbb{R}^{o}$ which do not match the system states ${\mathbf{x}}\in\mathbb{R}^{n}$. In the partial measurement scenario, with $o<n$, time-delay embedding techniques [10] can be adopted to recover hidden state components that might be fundamental to describe the dynamics [11]. Hereon we provide a description of this strategy – see Fig. 2 for an overall sketch.
First, time-delay embedding method consists in lifting the low-dimensional observations $y(t)$ (considering $o=1$ for simplicity) into high-dimensional space by stacking time-shifted copies of the measurements $y(t)$ by means of vectors $\mathbf{a}(t;w,\zeta)=[y(t),y(t+\zeta),y(t+2\zeta),\ldots,y(t+(w-1)\zeta)]^{\top}\in\mathbb{R}^{w}$, where $w$ and $\zeta$ indicate the number of delay embeddings and the lag time, respectively. The resulting matrix $\mathbf{A}$ is termed as Hankel matrix and has the following form (taking $\zeta=1$):
$\mathbf{A}=\begin{bmatrix}y(t_{1})&y(t_{2})&\cdots&y(t_{T-w})\\ y(t_{2})&y(t_{3})&\cdots&y(t_{T-w+1})\\ \vdots&\vdots&\ddots&\vdots\\ y(t_{w})&y(t_{w+1})&\cdots&y(t_{T})\end{bmatrix}=\begin{bmatrix}\mathbf{a}(t_{1}),\mathbf{a}(t_{2}),\ldots,\mathbf{a}(T-w)\end{bmatrix}\in\mathbb{R}^{w\times(T-w)}.$ (16)
Second, a Singular Value Decomposition (SVD) $\mathbf{A}=\mathbf{U}\mathbf{S}\mathbf{V}^{\top}$ is performed to extract time-delay coordinates approximating the Koopman operator in finite dimension [40], thereby providing a linearised description of the system dominant dynamics [27]. The matrices $\mathbf{U}\in\mathbb{R}^{w\times w}$ and $\mathbf{V}\in\mathbb{R}^{(T-w)\times(T-w)}$ are two orthogonal matrices whose columns are termed left and right singular vectors, respectively; ${\mathbf{S}}\in\mathbb{R}^{w\times(T-w)}$ is a pseudo-diagonal matrix collecting the associated singular values. The reference system formed by the columns of $\mathbf{U}$ is diffeomorphic to the original attractor of the system dynamics, under the conditions given by the Takens’ embedding theorem [41]. The number of time-delay coordinates often can be significantly reduced to $\eta\ll w$ and can be chosen according to the decay of the corresponding singular values, thus resulting in a truncated SVD: $\mathbf{A}\approx\tilde{\mathbf{U}}\tilde{\mathbf{S}}\tilde{\mathbf{V}}^{\top}$, with $\tilde{\mathbf{U}}\in\mathbb{R}^{w\times\eta},\tilde{\mathbf{S}}\in\mathbb{R}^{\eta\times\eta},\tilde{\mathbf{V}}\in\mathbb{R}^{(T-w)\times\eta}$. The selected $\eta$ coordinates are expected to retrieve the unobserved, hidden, state variables, and therefore can be leveraged by SINDy in the offline training phase, to better capture the dynamics and to improve prediction performance [11]. Ideally $\eta=n$, being $n$ the state dimensionality of the partially-observed system under consideration.
An additional advantage of time-delay embedding is that it allows the reconstruction of the original variables through back-projection. This provides an explicit expression for the function $\mathbf{h}$ that maps the identified latent states to the corresponding observation vector and its Jacobian $\mathbf{H}$. This enables the online prediction phase to be performed in the time-delay embedding coordinates and to exploit low dimensional observations in the correction phase, thus maintaining the EKF-SINDy online procedure unchanged and avoiding the computational cost of computing the Jacobian (8).
The performance of the time-delay embedding could depend significantly on the parameters choices, such as the number of embeddings $w$ and the time lag $\zeta$. As general and qualitative guidance, time delay parameters should be chosen to allow for the attractor to properly unfold, thus occupying as much of the embedded phase space as possible [11]. More quantitative techniques, based on relation or mutual information [42], could be employed as in [43, 44].
Figure 2: Time-delay emebdding procedure. Over the offline phase, time-series observations of the system are time-delay embedded to form the Hankel matrix $\mathbf{A}$. The time-delay coordinates are then obtained by projecting the delayed signals onto the dominant $\eta$ left-singular values of the SVD decomposition (ı.e., columns of $\tilde{\mathbf{U}}$) and rescaled by dividing them by the corresponding singular values, stored in $\tilde{\mathbf{S}}$. Finally, SINDy is trained offline on these time-delay variables, approximating the full state space variables of the system under consideration (we stressed this aspect by indicating $\eta=n$, even though in general $\eta$ could be different from $n$). Once trained offline, SINDy could be employed in the online prediction phase to forecast the evolution of the time-delay variables, from which the observations can be reconstructed by back-projection (ı.e. by multiplication for $\tilde{\mathbf{U}}\tilde{\mathbf{S}}$). Then the reconstructed observations can be simply extracted by looking at the first column, which contains the unshifted reconstructed signal, and compared to the actual testing observations (assimilated online) to perform the final correction phase. It follows that state-to-observation map, $\mathbf{h}$, is straightforwardly obtained as $\mathbf{h}=\mathbf{e}_{1}^{\top}\tilde{\mathbf{U}}\tilde{\mathbf{S}}$, where $\mathbf{e}_{1}$ is the basis vector extracting the first column of $\tilde{\mathbf{U}}\tilde{\mathbf{S}}$.
2.3 Joint estimation of non-autonomous dynamical systems
@@
The goal of joint estimation is to simultaneously estimate $\mathbf{x}$ and the parameters $\boldsymbol{\phi}$ [45]. To reach the scope, we increase the number of state variables by augmenting the state vector as $\boldsymbol{\varkappa}=[\mathbf{x}^{\top},\boldsymbol{\phi}^{\top}]^{\top}$. We consider the case in which it is impossible to associate a dynamic description to the evolution of $\boldsymbol{\phi}$. In this setting, the prediction stage of $\boldsymbol{\phi}$ is performed by a random walk driven by a process noise $\mathbf{q}_{\phi}$ as in the following:
@@
where $\mathbf{q}_{\phi}$ is assumed to be sampled from a white, zero mean stochastic process $\mathbf{q}_{\phi}\sim\mathcal{N}(\mathbf{0},\mathbf{Q}_{\phi})$ with a diagonal covariance matrix $\mathbf{Q}_{\phi}\in\mathbb{R}^{l\times l}$. Keeping the definition of $\mathbf{q}_{x}$ and $\mathbf{Q}_{x}$, we introduce a unique process noise vector $\mathbf{q}=[\mathbf{q}_{x},\mathbf{q}_{\phi}]^{\top}$ and a unique diagonal covariance matrix $\mathbf{Q}$ collecting the entries of $\mathbf{Q}_{x}$ and $\mathbf{Q}_{\phi}$. Finally, we update the definition of $\mathbf{f}:\mathbb{R}^{n+l}\rightarrow\mathbb{R}^{n+l}$ to include both the function modelling the system dynamics and the parameter evolution ruled by Eq. (17).
@@
Also in case of joint estimation, it is possible – and convenient – to rely on SINDy. Specifically, the dependence of system dynamics on $\boldsymbol{\phi}$ can be accounted for by stacking snapshots $\boldsymbol{\varkappa}$ accounting for $I$ different realisations of $\boldsymbol{\phi}$, thereby obtaining a modified version $\mathbf{X}_{\varkappa}$ of the snapshot matrix. Thus, SINDy is employed to reconstruct the functional relation between $\mathbf{X}_{\varkappa}$ and $\dot{\mathbf{X}}$. It is unnecessary to modify the definition of $\dot{\mathbf{X}}$, as we have assumed to drive the dynamics of $\boldsymbol{\phi}$ with a random walk. It is worth noticing that also the definition function library $\boldsymbol{\Theta}_{f}(\boldsymbol{\varkappa})=[\theta_{1}(\boldsymbol{\varkappa}),\dots,\theta_{p}(\boldsymbol{\varkappa})]\in\mathbb{R}^{p}$ must be updated to account for the dependence on $\boldsymbol{\varkappa}$. Specifically, the matrix $\boldsymbol{\Theta}_{f}(\mathbf{X}_{\varkappa})$, defined in Eq. (10), is modified by applying $\boldsymbol{\Theta}$ to the rows of $\mathbf{X}_{\varkappa}$ obtaining:
$\boldsymbol{\Theta}_{f}(\mathbf{X})=\begin{bmatrix}|&|&|&|&&|&|&\\ 1&\mathbf{X}_{\varkappa}&\mathbf{X}_{\varkappa}^{P_{2}}&\mathbf{X}_{\varkappa}^{P_{3}}&\cdots&\text{sin}(\mathbf{X}_{\varkappa})&\text{cos}(\mathbf{X}_{\varkappa})&\cdots\\ |&|&|&|&&|&|&\end{bmatrix}.$ (18)
Looking at the quadratic nonlinearities, Eq. (11) is updated to include the dependence on $\boldsymbol{\phi}$ as follows:
$\mathbf{X}^{P_{2}}_{\varkappa}=\begin{bmatrix}x^{2}_{1}(t_{1})&x_{1}(t_{1})x_{2}(t_{1})&\cdots&x_{1}(t_{1})\phi_{1}(t_{1})&\cdots&x^{2}_{n}(t_{1})&\cdots&{\phi}^{2}_{l}(t_{1})\\ x^{2}_{1}(t_{2})&x_{1}(t_{2})x_{2}(t_{2})&\cdots&x_{1}(t_{2})\phi_{1}(t_{2})&\cdots&x^{2}_{n}(t_{2})&\cdots&{\phi}^{2}_{l}(t_{2})\\ \vdots&\vdots&\ddots&\vdots&\ddots&\vdots&\ddots&\vdots\\ x^{2}_{1}(t_{T})&x_{1}(t_{T})x_{2}(t_{T})&\cdots&x_{1}(t_{T})\phi_{1}(t_{T})&\cdots&x^{2}_{n}(t_{T})&\cdots&{\phi}^{2}_{l}(t_{T})\end{bmatrix}.$ (19)
Higher order linearities and trigonometric functions can be similarly treated.
Notably, setting $\dot{\boldsymbol{\phi}}=0$ implies that the parameter values can be updated only in the correction phase of the EKF. This is allowed by updating Eq. (15a) to define the Jacobian $\mathbf{F}\in\mathbb{R}^{(n+l)\times(n+l)}$ including the dependence of system dynamics on $\boldsymbol{\phi}$ as in the following:
[…]
where the rows of $\mathbf{F}$ containing $\mathbf{0}$ entries, with $\mathbf{0}$ representing null matrices of suitable dimensions, are associated with the random walk equations. When used to model $\mathbf{h}$ for joint estimation, SINDy instead reconstructs the mapping between $\mathbf{X}_{\varkappa}$ and $\mathbf{Y}$. Also the definition of $\mathbf{Y}$ must not be modified as system parameters can not, in general, be observed. We highlight that the procedure can not explicitly handle the dependence on other parameters not included in $\boldsymbol{\phi}$, for which it is impossible to provide a separate estimate. This issue will be addressed in a future extension of the current work.
@@
$\dot{\mathbf{x}}(t)$ $=\mathbf{f}(\mathbf{x}(t),\mathbf{b}(t)),$ (21a)
@@
Similar modifications apply to Eqs. (3a) and (3b). The other equations of the filter remain unchanged if we assume to know $\mathbf{b}$ that is in the case of an exogenous input. This is true, for instance, if we consider a seismic excitation thanks to the presence of monitoring networks [46]. If $\mathbf{b}$ was unknown, it would be possible to modify the formulation of the filter to perform input–state–parameter estimation [47, 48, 49, 50]. Also this further extension of the method will be considered in future.
The external forcing term can be handled via SINDy by constructing a snapshot matrix $\mathbf{X}_{\varkappa b}$ assembling different snapshots $[\boldsymbol{\varkappa}(t_{j}),\mathbf{b}(t_{j})]^{T}$ determined for several realisations of $\boldsymbol{\phi}$. As a result, SINDy is employed to reconstruct the functional relation between $\mathbf{X}_{\varkappa b}$ and $\dot{\mathbf{X}}$ (and between $\mathbf{X}_{\varkappa b}$ and $\mathbf{Y}$ when used to model $\mathbf{h}$). Once again, the function library $\boldsymbol{\Theta}_{f}$ should be updated to account for the dependence on $\mathbf{b}$, similarly to what has been previously done in Eqs. (10) and (11) to consider the dependence on $\boldsymbol{\phi}$.
The complete procedure for the joint estimation of non-autonomous dynamical system are reported in Algorithms 1 and 2. In the algorithms, we consider the most general case in which both $\mathbf{f}$ and $\mathbf{h}$ are identified by using SINDy. The training of the SINDy model, and therefore the assembling of the snapshot matrices, must be done in an offline phase a priori with respect to the application of the procedure. Algorithm 1 is devoted to the illustration of this phase assuming to know $\mathbf{f}$ and $\mathbf{h}$. This case is of interest since it will be considered in the results section. As previously remarked, it is also possible to estimate $\mathbf{f}$ and $\mathbf{h}$ using noise tolerant versions of SINDy. During the online phase, the procedure is applied to update the estimates of the system by using incoming data. Algorithm 2 refers to this second part.
@@
12:Compute $\boldsymbol{\xi}^{f}_{i}=\underset{\boldsymbol{\xi}^{\prime}_{i}}{\arg\min}\|\dot{\mathbf{X}}_{i}-\boldsymbol{\Theta}_{f}(\mathbf{X}_{\varkappa b})\boldsymbol{\xi}^{\prime}_{i}\|_{2}+\mathcal{L}(\boldsymbol{\xi}^{\prime}_{i})$ for $i=1,\ldots,n$
13:Compute $\boldsymbol{\xi}^{h}_{i}=\underset{\boldsymbol{\xi}^{\prime}_{i}}{\arg\min}\|{\mathbf{Y}}_{i}-\boldsymbol{\Theta}_{h}(\mathbf{X}_{\varkappa b})\boldsymbol{\xi}^{\prime}_{i}\|_{2}+\mathcal{L}(\boldsymbol{\xi}^{\prime}_{i})$ for $i=1,\ldots,o$
@@
Algorithm 1 EKF-SINDy joint estimation ($\boldsymbol{\varkappa}=[\mathbf{x},\boldsymbol{\phi}]^{\top}$) of non-autonomous dynamical systems: offline phase ($\mathbf{f}$ and $\mathbf{h}$ available)